7. Find the area under the curve y = x² +1 on the interval (1,2). a. 73 b.3 d. 4 e. 7
Answer #1

let A be the area under the come y=x²+1 an the intonal [1,2] Then . 2 A= (y dx = 5x + 1) dx = x + x 1² - - ( 22)-bit 1921- +let A be the area under the come y=x²+1 an the intonal [1,2] Then . 2 A= (y dx = 5x + 1) dx = x + x 1² - - ( 22)-bit 1921- +